OpenStudy (anonymous):

Can someone please help and explain this to me?? Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 4x2 + 6x + 2 = 0

OpenStudy (anonymous):

B squared minus 4ac

OpenStudy (anonymous):

Here

OpenStudy (mathstudent55):

The part b^2 - 2ac is called the discriminant. If the discriminant is positive, the equation has 2 real solutions. If the discriminant equals zero, there is one real solution. If the discriminant is negative, there are no real solutions.

Nnesha (nnesha):

i guess you meant 4ac ? right ?....^

OpenStudy (mathstudent55):

LOL, yes, correct.

OpenStudy (jacob902):

Solving quadratic equations using quadratic formula? Question 1 (Multiple Choice Worth 1 points) What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth? x ≈ −3.23 and x ≈ 0.23 x ≈ −2.72 and x ≈ −0.28 x ≈ 0.28 and x ≈ 2.72<----------- x ≈ −0.23 and x ≈ 3.23 For this one i keep getting 2.73 and 0.28 but the answer isnt there soo i just picked can you tell me if im right? Question 2 (Multiple Choice Worth 1 points) How can you tell when a quadratic equation has no real solutions? when the radicand is negative<---------- when b in the quadratic formula is greater than the radicand when the radicand equals zero when the radicand is not a perfect square For this one im almost positive its the answer i have the arrow pointed 2 but just want you to clarify it Question 3 (Multiple Choice Worth 1 points) For which value of x does the graph of y = 3x2 − 2x − 5 cross the x-axis? -5/3 −1 −2 -3/5 For this one i just need help i have noo freaking clue how to do it... Question 4 (Fill-In-The-Blank Worth 5 points) Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Numerical Answers Expected! Answer for Blank 1: Again this one i have no clue... Question 6 (Multiple Choice Worth 1 points) What are the exact solutions of x^2 = 4 −7x I thought it was x=7x+-the square root of 33 divided by 2 i think im right but im not sure Anyone please help the major questions i need help with is 4 and 3 but please try to clarify if the one i think i know are right are right Thank you :) Follow 2 answers Report Abuse Answers Krishnamurthy Best Answer: 1). 4x^2 + 3 = -12x 4x^2 + 12x - 3 = 0 x = (-12 +/- sqrt 96)/8 x = (-3 +/- sqrt 6)/2 x = -0.275255 x = 2.72474 Answer: x ≈ −2.72 and x ≈ −0.28 2). How can you tell when a quadratic equation has no real solutions? when the radicand is negative 3). For which value of x does the graph of y = 3x^2 - 2x - 5 cross the x-axis? -1 4). Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. b^2 - 4ac 5). Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0 b^2 - 4ac = 49 - 24 = 25 6). x^2 = 4 - 7x x^2 + 7x - 4 = 0 x = (-7 +/- sqrt 65)/2 Krishnamurthy · 3 years ago 0 Thumbs up 0 Thumbs down Comment Report Abuse Jeff Aaron 1. 4x^2 + 3 = -12x 4x^2 + 12x + 3 = 0 x = (-12 +/- sqrt(12^2 - 4(4)(3))) / (2*4) x = (-12 +/- sqrt(144 - 48)) / 8 x = (-12 +/- sqrt(96)) / 8 x = (-12 +/- sqrt(96)) / 8 x = -1.5 +/- sqrt(1.5) x =~ -0.28 or -2.72 2. How can you tell when a quadratic equation has no real solutions? When the discriminant is negative (assuming the coefficients are real). 3. 3x^2 - 2x - 5 = 0 x = (-(-2) +/- sqrt((-2)^2 - 4(3)(-5))) / (2*3) x = (2 +/- sqrt(4 + 60)) / 6 x = (2 +/- sqrt(64)) / 6 x = (2 +/- 8) / 6 x = (10 or -6) / 6 x = 5/3 or -1 6. x^2 = 4 - 7x x^2 + 7x - 4 = 0 x = (-7 +/- sqrt(7^2 - 4(1)(-4))) / (2*1) x = (-7 +/- sqrt(49 + 16)) / 2 x = (-7 +/- sqrt(65)) / 2 x = -3.5 +/- sqrt(16.25)